What is the pH of a soft drink in whoch the major buffer ingredients are 6.3 g of NaH2PO4 and 8.0 g of Na2HPO4 per 355 mL of solution?
Answer #1

mixture of NaH2PO4 and Na2HPO4 act as acidic buffer

pH = pKa + log [Na2HPO4] / [NaH2PO4]

[NaH2PO4] = 7.21

[NaH2PO4] = (6.3 / 119.98)(1000 / 355)

[NaH2PO4] = 0.148 M

[Na2HPO4] = (8.0 / 141.96) (1000 / 355)

[Na2HPO4] = 0.159 M

pH = 7.21 + log [0.159] / [0.148]

pH = 7.24